A sound wave of frequency 300Hz travels with speed 340 m/s along the negative x-direction . Each point of the medium moves up and down through a total distance of 5.0 mm. Determine the necessary parameters to represent the wave mathematically.

1
Expert's answer
2022-03-09T15:55:08-0500

Explanations and Calculations.


"\qquad\qquad
\begin{aligned}
\small \lambda&=\small \frac{v}{f}=\frac{340}{300}=1.13\,m\\
\small k&=\small \frac{2\pi}{\lambda}=\frac{30\pi}{17}=5.54\,m^{-1}\\
\small \omega&=\small 2\pi f= 600\pi=1885\, rads^{-1}\\
\\
\small y(x,t)&=\small A\sin(kx+\omega t)\\
&=\small \frac{5}{2}\sin(5.54x+1885t)\\
&=\small 2.5\sin(5.54x+1885t) 
\end{aligned}"